FGC Software Design Review - LHC commissioning_1_ by hcj

VIEWS: 7 PAGES: 12

									        Wednesday 13th February - last 24 hours
   n    07:41 Programmed dump of fill 3556 (510 bunches) after 5h 26min
   n    09:01 Start injection of fill 3557 (1347 bunches)
   n    10:30 Stable beams, initial lumi ~90E30
   n    11:32 Beam dump
         ¨ BLM spurious trigger on a BLM optical board in R1
   n 14:07 Stable beams fill 3558, initial lumi above 100E30
   n 15:15 Beam dump
         ¨ RF fault
   n 16:25 LHC in access mode (RF + MQK)
   n 20:25 Patrol started in PM45 (lost during access)
   n 21:39 Ready for injection
         ¨ dump test for XPOC data collection test
   n 22:40 Stable beams fill 3559 – initial lumi ~120E30
   n 07:00 Beams dumped by OP
         ¨ ~2 pb-1 in 8h30m                                    Thanks to Matteo Solfaroli
                                                                    for the slides!

13-02-2013                                  LHC status                                      1
        Total integrated luminosity at 1.38 TeV




                  3556                                             3559
                 (500b)
                          3557    3558




                                              ATLAS total now ~3.5 pb-1
13-02-2013                       LHC status                               2
        Last 24 hours




             3557   3558                3559


                                           Abort gap
                                         high cleaning
                                             effect




13-02-2013                 LHC status                    3
        Fill 3557


                                                 Intensity ~ 1.2E11 p/bunch
                                                    Injection B1 then B2
                                           Big losses on B1 at start of collisions



                                                     Peak lumi ~95E30



                    Improvement on lifetime
                    (~5h) when trimming the
                    octupoles down from 3.0
                    to 2.6 units in collisions




13-02-2013                         LHC status                                    4
        Fill 3557
             Strong excitation visible on B1 and B2 around 0.29 (8 kHz line?) which
         becomes more pronounced when going into collision. The line is visible already
                 before and seems to be most pronounced in the vertical plane.




                                  B1H                                     B1V
                      0.29                              0.29




                                  B2H                                     B2V
                    0.29                                 0.29
13-02-2013                                 LHC status                                     5
        Fill 3558
                     Lifetime dip and losses improved
                          changes wrt previous fill
             • Some time spent at flat-top to wait for the tunes to stabilise
             • Alternating injection B1 and B2
             • Reduced landau octupoles strength on the second part of the
               collision BP from 3 to 2.6 (~120 A to 100A)




                                                            Peak lumi ~140E30
13-02-2013                                 LHC status                           6
        Fill 3558 - dump

              At 15:15 beam dump due to trip of RF module M1B1

                        Alarms on RF:
   ACSModuleM1B1 : Crowbar Grid current (LV side) out of
   range. Else HV-RF specialist intervention required to
   check and in case replace crowbar.


   ACSLine4B1: Main HV Power not at Nominal
   ACSLine1B1: Main HV Power not at Nominal
   ACSLine3B1: Main HV Power not at Nominal
   ACSLine2B1: Main HV Power not at Nominal
   -> The power converter is not delivering a minimal
   nominal voltage (set to 40kV). Power converter off or
   wrong voltage output. Measurement system faulty. Check
   power converter settings and acquisition with expert
   interface.
   In case call power converter piquet or access required.


13-02-2013                              LHC status               7
        Access – Tuesday evening




             n Access was given to RF zone to fix the fault on M1B1
             n In parallel access for the MQK – a test was performed to
               check the MKQ functionality
             n During the access the patrol on PM45 sector 1 was lost – had
               to be redone




13-02-2013                              LHC status                            8
        Fill 3559
                                                               Peak lumi ~155E30


   Octupoles strength further lowered while
    in stable beams down to 1 unit, then
             trimmed back to 1.2

         Big improvement on lifetime



         Octupoles to 2                            AGC had to be switched ON continuously during the fill
             units


                  Abort gap    Octupoles to
                 cleaning ON    1.2 units




13-02-2013                                    LHC status                                               9
        Plans

     n Refill for physics
     n (beam unavailable from Booster this morning for 4
       hours)
     n This afternoon – VdM – INDIVs
             ¨ Plus loss map
     n Operational development
     n Fill overnight to finish – RPs in to 13 sigma
     n 06:00 Thursday – start quench test program
             ¨ No ions




13-02-2013                      LHC status                 10
        Roman pots - request
     n At the end of the upcoming 1400 bunch fill we briefly
       insert the pots to the validated position of 22 sigma, just
       as a necessary condition for proceeding further in the
       plan. I remind that last night's insertion to
             ¨ 22 sigma from the 500 bunch fill was without any problem.
     n At the end of a low-intensity fill, possibly the VdM fill, we
       insert the pots to 13 sigma (where we would have
       physics acceptance) and do a quick loss map at
       dumping time.
     n If this loss map is ok, we insert to 13 sigma in the
       remaining fills.



                                                       Mario Deile
13-02-2013                              LHC status                         11
        Accesses

     n Accesses for quench test:
             ¨ Scope installation in IP6 (12L6). 2 hours. It has priority over Q6L8 scope
               installation
             ¨ Fast measurement system on RQ6.L8 (30 min in the UA83 by Jaromir
               Ludwig). Ramp to 300 A afterwards.
     n Access for circuit tests – in the shadow
             ¨ Change hardware limits on all Individual Powered Quads
                n EPC, 2 hours during working hours




13-02-2013                                  LHC status                                      12

								
To top